How does Young`s modulus change with rise in temperature? A load of 2kg produces an extension of 1mm in wire of 3mm in length and 1mm in diameter. Calculate the Young`s modulus of elasticity of the wire?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
stress =  Force /Area of contact => 20/(pi*0.5*0.5) N/mm^2
strain = delta(l)/l = > 1/3
Young’s mudulus = stress/strain
Hence,
Y =  20*3/(pi*0.5*0.5) = > 76.43312    N/mm^2